If I had answered 3/4 of my math test and it contained 75 questions, how many did I answer correctly?
Found 3 solutions by stanbon, Theo, richwmiller:
Answer by stanbon(75887)

If I Had answered 3/4 of my math test and it contained 75 questions how many did I answer correct.
---
You answered (3/4)75 = 56+ questions.
How many you answered correctly is still a mystery.
---
Cheers,
Stan H.

Answer by Theo(13342)

(3/4) * 75 = 56.25

you answered 56.25 correctly.

56.25/75 = .75

.75 * 1 = .75 * 4/4 = (.75*4)/4 = 3/4





Answer by richwmiller(17219)

If we assume that you meant to say that you got 3/4 of the answers correct then 3/4(75)= the number that you answered correctly.
BUT your problem doesn't say that.
It actually says that you didn't even answer 1/4 of the questions and that you put something for an answer for 3/4 of them.
3/4(75) is 56.25
How you manage to answer such a strange number of questions is beyond me.
Maybe you started an answer and didn't finish it.
But for you to get that strange number correct is impossible. Maybe you get partial credit.